OpenStudy (anonymous):

find the formula for the inverse f(x)=x+4

OpenStudy (anonymous):

An inverse of a function is itself, a function revolved around the line y=x. To find the inverse (typically) you can simply swap x and y. So you have: y=x+4 Switch x and y Giving: x=y+4 Now solve for y x-4=y Is the inverse.